OpenStudy (anonymous):

Simplify Algebraic Expression n^2 + n + 5 - (2n^2+n+1) just tried to do it by myself, not sure i got the right answer 1. I dropped the brackets 2. Collected Like Terms (n^2 - 2n^2), (n+n), (5+1) 3. n^2 - 2n^2 = -n^2 ? 4. n + n = 2n ? 5. 5 + 1 = 6 Answer = -n^2 + 2n + 6 ?? I'm pretty sure i'm wrong tho.

OpenStudy (anonymous):

\[n^2 + n + 5 - (2n^2+n+1) \rightarrow n^2 + n + 5 - 2n^2-n-1\] \[=n^2 - 2n^2+ n -n+ 5 -1\] \[=- n^2+ 0+ 4\] \[=- n^2 + 4\] is the required answer..

OpenStudy (anonymous):

@pufss

OpenStudy (anonymous):

The negative beside the bracket meant that i had to subtract all the stuff inside the parenthesis and not add?

OpenStudy (anonymous):

@dpasingh

OpenStudy (cggurumanjunath):